# FLoRA: An Advanced AI-Powered Engine to Facilitate Hybrid Human-AI Regulated Learning

### v2

@@ 3.2 From Writing Analytics to Personalised Feedback
The second function, the Academic Writing module (Figure 5), supports learners in aligning their language with academic conventions. This module draws on a back-end Map data structure and NLP phrase-matching to efficiently identify issues with tone, sentence complexity, and vocabulary usage. By delivering context-specific guidance grounded in academic expectations, the tool promotes students’ ability to plan and adapt their writing style to disciplinary norms, thus supporting SRL strategies related to goal-setting, strategic planning, and self-reflection
@@
The Originality function (Figure 6) evaluates the uniqueness of student text by flagging any instance where more than seven consecutive words match the background reading material, leveraging an n-gram algorithm for detection. Beyond supporting academic integrity, the explicit highlighting of copied segments prompts students to monitor their use of sources, reconsider how they synthesize information, and actively paraphrase in their own words. This encourages the SRL processes of self-evaluation and strategic revision, empowering learners to internalize standards for originality.
@@ 3.3 Chatbot Support with Multiple GenAI-powered Agents
With the rapid advancement of GenAI, its integration into educational contexts has become increasingly prevalent. GenAI-powered tools, such as chatbots, provide students with immediate access to information and personalized guidance, thereby offering significant potential to foster SRL. By facilitating processes such as setting learning objectives, selecting effective strategies, and reflecting on outcomes, GenAI tools can scaffold key stages of SRL. Within the FLoRA Engine, we have developed and integrated GenAI-based chatbot tools that leverage comprehensive learning contexts—including students’ prior knowledge, learning goals, and trace data logs—to systematically monitor progress and deliver real-time, actionable feedback. This continuous, contextualized support promotes metacognitive development, encouraging learners to reflect on their approaches and make adaptive adjustments throughout the learning process.
@@
To accommodate diverse educational scenarios, the chatbot tool can be configured and used in flexible ways. The Figure 8 employs a single chat window that supports multiple GenAI agents, each configured with distinct pre-prompts to specialise in specific domains or tasks. All agents share the conversation history, enabling learners to interact with a panel of experts within a cohesive dialogue. This collaborative environment mirrors authentic problem-solving situations, where engaging with multiple perspectives enhances the depth and breadth of feedback. By receiving collective input from specialized agents, students are better able to monitor their understanding, seek clarification, and receive nuanced guidance—thereby strengthening SRL components such as self-monitoring, strategy selection, and adaptive regulation.
@@
The Figure 9 adopts a multi-window configuration, assigning each GenAI agent to a separate, independent chat interface. In this setup, conversation histories are not shared, allowing each interaction to remain focused and distinct. This design is particularly effective in simulations or role-based scenarios where compartmentalized, parallel discussions are required. For example, in a negotiation exercise, one agent may serve as the client and another as the supplier, enabling learners to engage fully with each perspective. This separation supports SRL by encouraging learners to set targeted objectives, monitor progress within specific contexts, and evaluate outcomes independently for each scenario.
@@ 3.4 Collaborative Writing with Adaptive Support
A key innovation of the FLoRA collaborative writing tool is the integration of a GenAI-powered chat interface. This embedded intelligent agent augments group communication by delivering context-sensitive feedback, answering queries, and generating relevant suggestions. The adaptive, personalized scaffolding provided by GenAI supports learners as they engage in real-time goal setting, monitor both their own and their peers’ contributions, and reflect on collaborative strategies. Consequently, students are empowered to exercise critical SRL skills such as self-monitoring, strategic revision, and adaptive planning, all within a supportive, interactive environment.
The fusion of AI and collaborative writing in FLoRA advances the concept of hybrid human–AI regulated learning. By mediating peer dialogue and offering individualized feedback, the GenAI agent fosters metacognitive development and the co-construction of knowledge. Learners actively regulate their learning by setting objectives, adjusting approaches, and evaluating outcomes, benefitting from immediate support tailored to both personal and group needs. This synergy enhances the depth and quality of collaboration, increases student engagement, and directly cultivates SRL competencies.
@@ 4 FLoRA Engine Implementation in Research Studies
To illustrate its practical utility, efficacy, and adaptability, this section presents representative case studies from diverse educational contexts. These examples demonstrate how the FLoRA Engine supports a wide array of learning objectives and tasks, showcasing its versatility and positive impact on learning processes. Through an examination of its application across scenarios, we offer evidence for the Engine’s capacity to advance SRL and HHAIRL in authentic environments.
@@ 4.1 Case 1 – Comparing Human, Rule-Based, and GenAI Feedback
Key Research Findings: This study yielded several key findings regarding learners’ interactions with GenAI-based chatbots, human experts, and writing analytics tools. First, learners who engaged with GenAI-based chatbots exhibited more direct and pragmatic help-seeking behaviors, often bypassing metacognitive stages such as diagnosing their difficulties or evaluating the support received. This resulted in a non-linear help-seeking process focused on obtaining specific information quickly through direct, knowledge-oriented questions. In contrast, learners interacting with human experts demonstrated a more linear approach, employing indirect and evaluative questioning to establish shared understanding and maintain social dynamics (Chen et al., 2025a; Cheng et al., 2025). This divergence in questioning styles mediated the relationship between support type and academic performance: learners who asked direct questions—typically supported by GenAI—showed greater improvements in essay revisions due to efficient knowledge acquisition (Cheng et al., 2025). However, reliance on GenAI sometimes reduced metacognitive engagement, as learners tended to delegate metacognitive monitoring to the chatbot, potentially limiting deep learning and critical reflection (Chen et al., 2025a). Second, GenAI-based chatbots and human experts embodied distinct values. The chatbot fostered social comfort and autonomy by minimizing social judgment, but was limited in responsiveness and raised privacy concerns. Conversely, human experts provided personalised and responsive support but occasionally constrained learner autonomy. Learners’ value priorities varied based on personal preferences and contexts, resulting in inherent trade-offs in their learning experiences (Shen et al., 2025). Third, while the GenAI chatbot often facilitated immediate performance gains—such as higher essay scores—these benefits did not translate into sustained intrinsic motivation or long-term knowledge gains compared to human tutoring or independent work. The convenience of GenAI support at times led to “metacognitive laziness,” with learners engaging less in planning and monitoring, both vital for self-regulated learning (SRL) (Authors, 2024). Fourth, the design of formative writing analytics tools significantly influenced learners’ self-assessment processes. Tools offering high-affordance feedback, such as clear grammatical corrections, supported accurate self-assessment and substantive improvement in writing. Conversely, low-affordance tools, including those integrating Bloom’s taxonomy classifications, often hindered effective self-reflection and revision due to lack of actionable guidance or potential for confusion (Tang et al., 2024; Forehand, 2010). Finally, although learners generally preferred human feedback, preferences were shaped by the type of GenAI tool used. High-performance, interactive chatbots like ChatGPT (see Section 3.3) fostered stronger preferences for AI-supported learning, whereas less interactive or rule-based tools (see Section 3.2 and Appendix A) led learners to favour human tutors for their greater personalization and interactivity (Le et al., 2025; Cheng et al., 2025).
@@ 4.2 Case 2 – Clinical Reasoning Training
Key Research Findings: The study assessed students’ clinical reasoning skills by analysing their history-taking behaviours during interactions with VSPs. Some behaviours showed strong correlations with reasoning metrics like diagnostic accuracy, history-taking scores, and clinical knowledge test results. This is because these behaviours reflected essential components of effective clinical reasoning: organizing questions in a logical sequence aids in systematic data collection; asking specific, pathophysiologically relevant questions demonstrates a deeper understanding of disease processes; and summarising and reorganising information indicates the ability to synthesise patient data. These behaviours enhance students’ ability to accurately diagnose and understand clinical scenarios, thereby improving their performance on reasoning metrics. By integrating GenAI-driven VSPs into the FLoRA platform, the study offered scalable and interactive opportunities for students to practice history-taking in a realistic, simulated environment. This approach enabled dynamic patient consultations and provided immediate, personalised feedback, thereby enhancing students’ clinical reasoning skills (Chen et al., 2025b).
Beyond generating valuable insights in medical diagnostics, this study also demonstrates the versatility of the FLoRA platform for a wide range of experiential learning tasks. Its capacity to simulate interactive scenarios suggests potential applications in leadership training and other domains requiring experiential learning. The results highlight that integrating FLoRA with GenAI technologies can facilitate skill development and deliver personalised learning experiences across diverse educational contexts.
@@ 4.3 Case 3 – advancing analytics-based adaptive SRL scaffolding with GenAI
This project investigated learners’ SRL skills in both K–12 and higher education contexts, with a focus on how learning analytics and AI can be leveraged to support robust SRL processes in the era of AI. A central innovation of this work is the integration of learning analytics with GenAI to enhance the adaptivity of SRL scaffolding. Prior research has identified the limitations of rule-based analytics, which often fail to accommodate the unique strategies and conditions of individual learners (Authors, 2025b). To overcome these limitations, this project utilised real-time learning analytics to inform GenAI models, enabling the generation of personalised and contextually relevant SRL support.
@@
Key Research Findings:
The design of this project has been implemented in empirical studies across both K-12 and higher education settings. In the higher education context, Authors (2025b) conducted randomised controlled trial compared the three groups as described above: CN, Po and PwC groups. Using Ordered Network Analysis (ONA), the results showed that learners in the PwC group demonstrated significantly more effective and metacognitive SRL patterns, such as frequent consultation of task instructions and rubrics, compared to the other groups. Furthermore, learner compliance (i.e., the level of adherence to the suggested processes) with the scaffolds varied, with compliers showing clear adherence to scaffold suggestions and more frequent metacognitive processes, especially in writing phases, while non-compliers exhibited more linear and less strategic learning patterns. These findings highlight the importance of incorporating both learners’ real-time SRL processes and evolving conditions in scaffold design and provide empirical evidence supporting the scalability and effectiveness of GenAI-powered adaptive scaffolding for enhancing SRL in computer-based learning environments. In another study, Authors (2025d) explored the relationship between feedback literacy and SRL processes among secondary school students across different countries. Through K-medoids clustering of self-reported feedback literacy, the study identified two distinct clusters: Proactive Feedback Engagers and Moderate Feedback Engagers. ONA revealed that Proactive Feedback Engagers initially employed less effective learning strategies but adapted significantly following scaffolded feedback, demonstrating more balanced and strategic regulation. In contrast, Moderate Feedback Engagers began with more strategic initial approaches but showed less adaptability and responsiveness to scaffolding suggestions. These findings emphasize the dynamic, context-specific nature of feedback literacy and its interplay with SRL, highlighting the need for adaptive scaffolding tailored to students’ feedback literacy profiles to enhance engagement and SRL processes.
@@ 5.1 Advancing Hybrid Human-AI Regulation in Education
However, the findings also highlight the risks of over-reliance on AI systems, echoing concerns raised in prior studies (Authors, 2024a). For instance, the case study of ESL learners using GenAI-based support for writing revealed instances of "metacognitive laziness," where learners offloaded planning and monitoring to AI systems. While this improved task-specific outcomes in the short term, it may hinder the long-term development of learners’ critical SRL abilities, a finding corroborated by research on the pitfalls of AI augmentation in learning contexts (Molenaar, 2022a).

### v3

@@ 3.2 From Writing Analytics to Personalised Feedback
The second function, the Academic Writing module (Figure 6), supports learners in aligning their language with academic conventions. This module draws on a back-end Key-Value based data structure and phrase-matching to efficiently identify issues with tone, sentence complexity, and vocabulary usage. By delivering context-specific guidance grounded in academic expectations, the tool promotes students’ ability to plan and adapt their writing style to disciplinary norms, thus supporting SRL strategies related to goal-setting, strategic planning, and self-reflection
@@
The Originality function (Figure 7) evaluates the uniqueness of student text by flagging any instance where more than seven consecutive words match the background reading material, leveraging an n-gram algorithm for detection. Beyond supporting academic integrity, the explicit highlighting of copied segments prompts students to monitor their use of sources, reconsider how they synthesise information, and actively paraphrase in their own words. This encourages the SRL processes of self-evaluation and strategic revision, empowering learners to internalise standards for originality.
@@ 3.3 Chatbot Support with Multiple GenAI-powered Agents
With the rapid advancement of GenAI, its integration into educational contexts has become increasingly prevalent. GenAI-powered tools, such as chatbots, provide students with immediate access to information and personalised guidance, thereby offering significant potential to foster SRL. By facilitating processes such as setting learning objectives, selecting effective strategies, and reflecting on outcomes, GenAI tools can scaffold key stages of SRL. Within the FLoRA Engine, we have developed and integrated GenAI-based chatbot tools that leverage comprehensive learning contexts—including students’ prior knowledge, learning goals, and trace data logs—to systematically monitor progress and deliver real-time, actionable feedback. This continuous, contextualised support promotes metacognitive development, encouraging learners to reflect on their approaches and make adaptive adjustments throughout the learning process.
@@
To accommodate diverse educational scenarios, the chatbot tool can be configured and used in flexible ways. The Figure 9 employs a single chat window that supports multiple GenAI agents, each configured with distinct pre-prompts to specialise in specific domains or tasks. All agents share the conversation history, enabling learners to interact with a panel of experts within a cohesive dialogue. This collaborative environment mirrors authentic problem-solving situations, where engaging with multiple perspectives enhances the depth and breadth of feedback. By receiving collective input from specialised agents, students are better able to monitor their understanding, seek clarification, and receive nuanced guidance—thereby strengthening SRL components such as self-monitoring, strategy selection, and adaptive regulation.
@@
The Figure 10 adopts a multi-window configuration, assigning each GenAI agent to a separate, independent chat interface. In this setup, conversation histories are not shared, allowing each interaction to remain focused and distinct. This design is particularly effective in simulations or role-based scenarios where compartmentalised, parallel discussions are required. For example, in a negotiation exercise, one agent may serve as the client and another as the supplier, enabling learners to engage fully with each perspective. This separation supports SRL by encouraging learners to set targeted objectives, monitor progress within specific contexts, and evaluate outcomes independently for each scenario.
@@ 3.4 Collaborative Writing with Adaptive Support
A key innovation of the FLoRA collaborative writing tool is the integration of a GenAI-powered chat interface. This embedded intelligent agent augments group communication by delivering context-sensitive feedback, answering queries, and generating relevant suggestions. The adaptive, personalised scaffolding provided by GenAI supports learners as they engage in real-time goal setting, monitor both their own and their peers’ contributions, and reflect on collaborative strategies. Consequently, students are empowered to exercise critical SRL skills such as self-monitoring, strategic revision, and adaptive planning, all within a supportive, interactive environment.
The fusion of AI and collaborative writing in FLoRA advances the concept of hybrid human–AI regulated learning. By mediating peer dialogue and offering individualised feedback, the GenAI agent fosters metacognitive development and the co-construction of knowledge. Learners actively regulate their learning by setting objectives, adjusting approaches, and evaluating outcomes, benefitting from immediate support tailored to both personal and group needs. This synergy enhances the depth and quality of collaboration, increases student engagement, and directly cultivates SRL competencies.
@@ 4 FLoRA Engine Implementation in Research Studies
To illustrate its practical utility, efficacy, and adaptability, this section presents representative case studies from diverse educational contexts. These examples demonstrate how the FLoRA Engine supports a wide array of learning objectives and tasks, showcasing its versatility and positive impact on learning processes. Through an examination of its application across scenarios, we offer evidence for the Engine’s capacity to advance SRL and HHAIRL in authentic environments. Because the system can be configured in different ways across contexts, the case studies adopt different research objectives, data structures, and analytical approaches. For this reason, statistical reporting is not fully homogeneous; instead, it is adapted to the methodological requirements of each case.
@@ 4.1 Case 1 – Comparing Human, Rule-Based, and GenAI Feedback
Key Research Findings: This study yielded several key findings regarding learners’ interactions with GenAI-based chatbots, human experts, and writing analytics tools. First, first-order Markov models (FOMM), used as a process mining technique, were used to analyse help-seeking stages and activities within each identified process. These stages included Diagnosing Question, Asking Help, Evaluating Help, and Processing Help. The process mining results showed that learners who engaged more with GenAI-based chatbots often bypassed Diagnosing Question or Evaluating Help which are related to metacognition (Chen et al., 2025a). The Mann-Whitney U test results indicated a significant difference between the two groups in the Asking Help stage ratio (Z=3.334,ES=1.288,p<0.001Z=-3.334,ES=1.288,p<0.001), with the AI Group showing a significantly higher ratio than that of HE Group. This is because GenAI-based chatbot provided instantly accessible resources, which naturally encourage more executive and direct help-seeking behaviours. This divergence in questioning behaviours mediated the relationship between support type and academic performance: learners who asked direct questions—typically supported by GenAI—showed greater improvements in essay revisions due to efficient knowledge acquisition (Cheng et al., 2025). However, reliance on GenAI sometimes reduced metacognitive engagement, as learners tended to delegate metacognitive monitoring to the chatbot, potentially limiting deep learning and critical reflection (Chen et al., 2025a). Second, after the experiment, 59 participants agreed to participate in the semi-structured interviews (33 in the AI group and 26 in the HE group, lasting 35 minutes). The interview transcription data was analysed using thematic analysis assisted by NVivo 12. The results showed that the chatbot fostered social comfort and autonomy by minimizing social judgment, but was limited in responsiveness and raised privacy concerns. Conversely, human experts provided personalised and responsive support but occasionally constrained learner autonomy. Learners’ value priorities varied based on personal preferences and contexts, resulting in inherent trade-offs in their learning experiences (Shen et al., 2025). Third, in another study, learners’ motivation was measured using the Intrinsic Motivation Inventory (McAuley et al., 1989; Torbergsen et al., 2023) in the post-task to compare intrinsic motivation across groups on four subscales: Interest/Enjoyment, Perceived Competence, Effort/Importance, Pressure/Tension. One-way ANOVAs followed by Tukey’s HSD were performed for each subscale. No significant between-group differences were observed for Interest/Enjoyment ($F=1.087,p=0.358,\eta^{2}=0.029$), Perceived Competence ($F=0.453,p=0.716,\eta^{2}=0.012$), Effort/Importance ($F=1.152,p=0.332,\eta^{2}=0.030$), or Pressure/Tension ($F=0.546,p=0.652,\eta^{2}=0.015$). Complementary analyses of SRL process frequencies and process mining indicated that the GenAI condition involved fewer metacognitive activities. Thus, although the GenAI chatbot often yielded immediate performance gains (e.g., higher essay scores), these gains were not accompanied by increases in intrinsic motivation and did not generalise to longer-term knowledge outcomes relative to human tutoring or independent study. The convenience of GenAI support at times led to "metacognitive laziness", with learners engaging less in planning and monitoring, both vital for SRL (Fan et al., 2025). Fourth, the design of formative writing analytics tools significantly influenced learners’ self-assessment processes. An epistemic network analysis (Shaffer et al., 2016) was conducted to examine the impact of different feedback tools on learners’ writing outcomes. The results demonstrated that tools offering high-affordance feedback, such as clear grammatical corrections, supported accurate self-assessment and substantive improvement in writing. Conversely, low-affordance tools, including those integrating Bloom’s taxonomy classifications, often hindered effective self-reflection and revision due to lack of actionable guidance or potential for confusion (Tang et al., 2024; Forehand, 2010). Finally, although learners generally preferred human feedback, preferences were shaped by the type of GenAI tool used. High-performance, interactive chatbots like ChatGPT (see Section 3.3) fostered stronger preferences for AI-supported learning, whereas less interactive or rule-based tools (see Section 3.2 and Appendix A) led learners to favour human tutors for their greater personalization and interactivity (Le et al., 2025; Cheng et al., 2025).
@@ 4.2 Case 2 – Clinical Reasoning Training
Key Research Findings: The study assessed students’ clinical reasoning skills by analysing their history-taking behaviours during interactions with VSPs. By analysing 1,030 consultation dialogues, a Pearson correlation analysis was conducted, and students’ dialogues were coded into 12 categories. Some behaviours showed strong correlations (r = 0.208–0.225, $\chi^{2}$=6.019, all P < .05) with reasoning metrics like diagnostic accuracy, history-taking scores, and clinical knowledge test results. This is because these behaviours reflect essential components of effective clinical reasoning: organizing questions in a logical sequence aids in systematic data collection; asking specific, pathophysiologically relevant questions demonstrates a deeper understanding of disease processes; and summarising and reorganising information indicates the ability to synthesise patient data. These behaviours enhance students’ ability to accurately diagnose and understand clinical scenarios, thereby improving their performance on reasoning metrics. By integrating GenAI-driven VSPs into the FLoRA platform, the study offered scalable and interactive opportunities for students to practice history-taking in a realistic, simulated environment. This approach enabled dynamic patient consultations and provided immediate, personalised feedback, thereby enhancing students’ clinical reasoning skills (Chen et al., 2025b). However, GenAI can be prone to generating hallucinations, resulting in the presentation of inaccurate or fabricated information (Qian et al., 2025). Consequently, the GenAI-based VSPs may sometimes provide students with misleading responses. This can hinder students’ diagnostic reasoning by introducing errors, and thus raises important concerns about the reliability of using GenAI in medical education.
Beyond generating valuable insights in medical diagnostics, this study also demonstrates the versatility of the FLoRA platform for a wide range of experiential learning tasks. Its capacity to simulate interactive scenarios suggests potential applications in leadership training and other domains requiring experiential learning. The results highlight that integrating FLoRA with GenAI technologies can facilitate skill development and deliver personalised learning experiences across diverse educational contexts. And, because of the personalised and adaptive support throughout the learning process, thereby the development of SRL skills can also be facilitated.
@@ 4.3 Case 3 – advancing analytics-based adaptive SRL scaffolding with GenAI
This project investigated learners’ SRL skills in both K–12 and higher education contexts, with a focus on how learning analytics and AI can be leveraged to support robust SRL processes in the era of AI. A central innovation of this work is the integration of learning analytics with GenAI to enhance the adaptivity of SRL scaffolding. Prior research has identified the limitations of rule-based analytics, which often fail to accommodate the unique strategies and conditions of individual learners (Li et al., 2025a). To overcome these limitations, this project utilised real-time learning analytics to inform GenAI models, enabling the generation of personalised and contextually relevant SRL support.
@@
Key Research Findings: The design of this project has been implemented in empirical studies across both K-12 and higher education settings. In the higher education context, Li et al. (2025a) conducted a randomised controlled trial compared the three groups as described above: CN, Po and PwC groups. Using Ordered Network Analysis (ONA), learners’ SRL processes—and how those processes transition among one another—onto a two-dimensional space. Through a means-rotation method, differences were maximised on the first dimension, and a Wilcoxon test was subsequently performed. The results indicated that learners in the PwC group demonstrated significantly more effective and metacognitive SRL behaviors, such as frequently consulting task instructions and rubrics, compared to the other groups (PwC vs. CN: $W$ = 1944, $n$ = 112 , $p$ = 0.0213, $r$ = 0.707; PwC vs. Po: $W$ = 1149, $n$ = 109, $p$ = 0.0478, $r$ = 0.863). Furthermore, the study employed Dynamic Time Warping (DTW) with k k-means clustering to assess learners’ compliance (i.e., the level of adherence to the suggested processes) with the SRL scaffolds, and differences between compliers and non-compliers were mapped onto the ONA space. The results revealed that learners’ compliance varied significantly: compliers demonstrated clear adherence to scaffold recommendations and engaged in more frequent metacognitive processes, particularly during writing phases. In contrast, non-compliers exhibited a more linear, less strategic learning trajectory ($W$ = 596, $n$ = 54, $p$ < 0.000, $r$ = 0.548). These findings highlight the importance of incorporating both learners’ real-time SRL processes and evolving conditions in scaffold design and provide empirical evidence supporting the scalability and effectiveness of GenAI-powered adaptive scaffolding for enhancing SRL in computer-based learning environments. By using this approach, our tool demonstrates the capability of delivering personalised support that adapts not only to learners’ real-time SRL processes but also to dynamic learning conditions that shift throughout the task. This adaptivity, in turn, promotes improved compliance and long-term impacts in facilitating effective self-regulation in the age of AI.
However, it is important to note that these findings did not reveal a direct effect on learners’ writing performance, as measured by their scores. This outcome may be attributed to the tools’ primary focus on supporting SRL activities, which are intended to foster long-term learning development rather than immediate performance gains. Such an interpretation aligns with prior research suggesting that SRL interventions typically do not produce immediate or significant improvements in essay performance (Yan et al., 2024). In another study, Jin et al. (2025) explored the relationship between feedback literacy and SRL processes among secondary school students across different countries. Through K-medoids clustering of self-reported feedback literacy, the study identified two distinct clusters: Proactive Feedback Engagers and Moderate Feedback Engagers. ONA revealed that Proactive Feedback Engagers initially employed less effective learning strategies but adapted significantly following scaffolded feedback, demonstrating more balanced and strategic regulation. In contrast, Moderate Feedback Engagers began with more strategic initial approaches but showed less adaptability and responsiveness to scaffolding suggestions. These findings emphasise the dynamic, context-specific nature of feedback literacy and its interplay with SRL, highlighting the need for adaptive scaffolding tailored to students’ feedback literacy profiles to enhance engagement and SRL processes.
@@ 5.1 Advancing Hybrid Human-AI Regulation in Education
However, the findings also highlight the risks of over-reliance on AI systems, echoing concerns raised in prior studies (Fan et al., 2025). For instance, the case study of ESL learners using GenAI-based support for writing revealed instances of "metacognitive laziness," where learners offloaded planning and monitoring to AI systems. While this improved task-specific outcomes in the short term, it may hinder the long-term development of learners’ critical SRL abilities, a finding corroborated by research on the pitfalls of AI augmentation in learning contexts (Molenaar, 2022a).